OpenStudy (anonymous):

how do you find the derivative of f(x)=e^x/x^2 ?

OpenStudy (zzr0ck3r):

\[\frac{e^x}{x^2}=e^xx^{-2}\] use the product rue

OpenStudy (anonymous):

will the answer be f'(x)= -2xe^x + e^x/x^2 ?

OpenStudy (isaiah.feynman):

Use quotient rule. take \[e^{x}\] to be f and \[x^{2}\] to be g. The quotient rule is this... \[\frac{ f^{'}g-f g^{'} }{ g^{2} }\]

OpenStudy (anonymous):

which one is right , -2xe^x + e^x/x^2 or x^2e^x +2xe^x all over x^4 ??

OpenStudy (isaiah.feynman):

Both of them are right, choose the one you like. :)

OpenStudy (anonymous):

thank you :)

OpenStudy (isaiah.feynman):

Yeah, I just solved the derivative using the two methods and I got the same answer.

OpenStudy (anonymous):

okay thank you very much :)

OpenStudy (isaiah.feynman):

No problem.